# Are Image Distributions Indistinguishable to Humans Indistinguishable to Classifiers?

### v1

@@ C.1 Results of CIFAR-10
C.2 Visualization of crops
As shown in Fig. 7, we present the visualization of crops mentioned in Sec. 7.1 and Sec. 7.2.
@@ C.2 Visualization of crops
C.3 Frequency analysis on the combination of EDM2-XS and EDM2-XXL
@@
Table 12: Classification accuracy of ResNet-50 on EDM2-XS and EDM2-XXL combinations after applying band-pass filters.
@@
C.4 User study
Fig. 10 shows a screenshot of the interaction interface used in our user study. There are eleven participants. We design three groups of experiments, in each of which the participants are asked to classify 32 pairs of images. The first set involved distinguishing generated images from real ones, where the real images are from ImageNet-256, and the generated images are from U-ViT-H/2. The second set involved classifying two diffusion models with similar performance, specifically DiT-XL/2 and U-ViT-H/2. The final set included images from EDM2-XS and EDM2-XXL, which utilize the same training methodology but differ in the number of parameters, leading to different FID. For the first set, participants were required to identify the real images. In the second set, they were tasked with identifying DiT images; since they lacked prior exposure to these images, we provided some reference images from DiT and U-ViT during the test. In the third experiment, participants were asked to determine which images were of better quality.

### v2

@@ Appendix B Computational cost
Appendix C Additional results
C.1 Additional results from generative adversarial network
As shown in Tab. 14, we conducted experiments using StyleGAN-XL (Sauer et al., 2022), a state-of-the-art GAN model trained on ImageNet-256. Our results demonstrate that, despite the use of a discriminator during GAN training, the classifier can still easily distinguish between real and generated images. We argue that this is because the discriminator is trained jointly with the generator. During training, the generated data seen by the discriminator comes from a continuously evolving distribution, as the generator improves with each iteration. However, when using a classifier to distinguish between real and generated distributions, the generated distribution remains fixed.
Real dataset Generative model FID Classifier Accuracy (%)
I-256 StyleGAN-XL (Sauer et al., 2022) 2.30 ResNet-50 99.69
ViT-S 99.95
ConvNeXt-T 96.85
C.2 Results of CIFAR-10
C.3 Self-supervised classifiers for Text-to-Image distribution classification
As shown in Tab. 15, we report the distribution classification accuracy of self-supervised classifiers in text-to-image scenarios. Notably, these classifiers achieve high accuracy in distinguishing between different text-to-image models. This suggests that there are significant differences between text-to-image generative models, allowing even self-supervised classifiers to easily distinguish them.
Table 15: Four-way distribution classification on text-to-image. Linear classifier on self-supervised features also achieves high accuracy in distinguishing four distributions: COCO (Lin et al., 2014), Pixart-$\alpha$ (Chen et al., 2023), SDXL (Podell et al., 2023), and Playground-v2.5 (Li et al., 2024), using only 5k or 10k training samples per dataset. We use ViT-B as the backbone.
Training samples Self-supervised method Accuracy (%)
5k MAE 87.77
5k MoCo v3 90.48
10k MAE 91.67
10k MoCo v3 90.90
C.4 Visualization of crops
As shown in Fig. 7, we present the visualization of crops mentioned in Sec. 7.1.
@@ C.4 Visualization of crops
C.5 Frequency analysis on the combination of EDM2-XS and EDM2-XXL
@@
Table 16: Classification accuracy of ResNet-50 on the combination of EDM2-XS and EDM2-XXL after applying band-pass filters.
@@
C.6 User study
Fig. 10 shows a screenshot of the interaction interface used in our user study. The study involved nineteen participants, and we designed three groups of experiments, each requiring participants to classify 32 pairs of images. In the first set, participants distinguished between generated and real images, with real images sourced from ImageNet-256 and generated images from U-ViT-H/2. In the second set, they classified images between two diffusion models with similar performance: DiT-XL/2 and U-ViT-H/2. For the final set, participants evaluated images from EDM2-XS and EDM2-XXL, which share the same training methodology but differ in parameter count, resulting in different FID scores and visual quality. In the first set, participants were asked to identify the real images. In the second, they were tasked with identifying DiT images, and reference images from DiT and U-ViT were provided during the test. In the third experiment, participants judged which images were of higher quality. All nineteen participants were graduate students with substantial experience in machine learning. They were allowed to zoom in on the images during the experiment, which was conducted on 27-inch 4K displays. All participants had corrected vision of 1.0 (standard normal vision), and their ages ranged from 22 to 26. Each participant completed the experiment within an hour and was compensated \$10.
@@ C.6 User study
Appendix D Binary classification as a measure of distribution distance
We employ a classifier $C(x)$ to distinguish between the real data distribution $p_{\text{data}}(x)$ and the generated data distribution $p_{g}(x)$. By training $C(x)$ using the binary cross-entropy loss:
$L(C)=-E_{x\sim p_{\text{data}}(x)}[\log(C(x))]-E_{x\sim p_{g}(x)}[\log(1-C(x))],$ (5)
The optimal classifier that minimizes this loss is:
$C^{*}(x)=\frac{p_{\text{data}}(x)}{p_{\text{data}}(x)+p_{g}(x)}.$ (6)
Substituting $C^{*}(x)$ back into the loss function yields:
$L(C^{*})=-\log(4)+2\,\text{JSD}(p_{\text{data}}(x)\|p_{g}(x)),$ (7)
where JSD denotes the Jensen-Shannon Divergence—a direct measure of the distance between the two distributions.
In our paper, we use classification accuracy to evaluate how well the classifier distinguishes between real and generated data because it provides an intuitive and interpretable metric. Although accuracy is non-differentiable and unsuitable for direct optimization, training the classifier with the binary cross-entropy loss—a convex surrogate—often leads to improved accuracy. This correlation suggests that accuracy can serve as a proxy for changes in the loss function, reflecting the distance between the generated and real data distributions.
(a) Accuracy vs. low-frequency filter threshold
(b) Accuracy vs. high-frequency filter threshold
Figure 11: Comparison of model accuracy across different filter thresholds using rectangular and circular filters.
Appendix E Relationship between distribution classification and FID
The Fréchet Inception Distance (FID) is a widely used metric for evaluating the quality of generative models. It relies on feature extraction networks trained on datasets such as ImageNet and assumes that the extracted feature vectors follow a multivariate Gaussian distribution. FID calculates the Fréchet distance between these Gaussian distributions to measure discrepancies between the real and generated data. Meanwhile, as noted by Kynkäänniemi et al. (2022), FID can decrease simply by aligning the histograms of top-N classifications, without necessarily improving the perceptual quality of the generated images. Additionally, recent works like Karras et al. (2023) and Tian et al. (2024) report FID scores close to those of the ImageNet validation set, suggesting limitations in FID’s sensitivity to certain distribution differences.
In contrast, our classifier-based approach offers a more direct measure of the distance between distributions without relying on the Gaussian assumptions inherent in FID. By training a classifier to distinguish between real and generated data, we obtain an intuitive and interpretable metric that reflects the actual distribution differences. This method complements commonly used metrics such as FID and Inception Score (IS), providing an alternative perspective on evaluating generative models.
Appendix F Comparison of rectangular and circular filters
Rectangular and circular filters are common techniques in image filtering. In this paper, we chose to implement a rectangular filter following the official FreeU implementation (Si et al., 2024), due to its simplicity and computational efficiency. For comparison, Fig. 11 presents initial results using a circular mask, denoted as U-H/2 and I-256 (Ideal). The results from both implementations are similar, and we chose to use the rectangular filter for its computational efficiency.